# Existence and Uniqueness of Solutions for Steady Triple-Deck Equations

### v1

@@ Abstract.
Triple-deck structures arise in boundary-layer flows when small localized wall perturbations induce moderate mean-flow distortion. Most existing studies in this field focus on the derivations of the asymptotic and numerical solutions of this system, while rigorous mathematical results remain scarce. In this paper, we establish, for the first time, the existence and uniqueness of the solutions to the steady triple-deck equations. The steady triple-deck equations can be reduced to an Airy equation with frequency as a parameter via Fourier transformation. The main difficulty is that the Airy equation degenerates at zero frequency, hindering uniform estimates of the solution with respect to the frequency parameter. A key contribution is that we find a subtle function involving the distribution of zeros of the Bessel function to overcome the degeneracy by analyzing the structure of the Airy function. Additionally, we derive a higher-frequency estimate and modify an elliptic multiplier to obtain uniform estimates of the solution.
@@ 1. Introduction
In 1904, Prandtl [21] proposed the boundary layer theory to address attached flow problems, achieving remarkable success in fluid dynamics research. However, when dealing with flow separation induced by adverse pressure gradients, this theory encounters singularity issues. Notably, Landau and Lifshitz [14] revealed that the velocity component normal to the body surface grows unboundedly (termed the Goldstein singularity), which hinders the solution from being further extended to the reverse-flow region [6]. To overcome these difficulties, scholars including A.F. Messiter [17] and K. Stewartson [27] developed the triple-deck theory from the late 1960s to the early 1970s. In this theory, three distinct layers—a viscous sublayer, a main layer, and an upper layer—appear in the transverse direction, for which uniformly valid approximate solutions are derived by the asymptotic matched method. For instance, Messiter successfully eliminated the Goldstein singularity in his study on the flow near the trailing edge of an incompressible flat plate, demonstrating the theory’s mathematical resolution to singular behaviors [17].
In physics, the triple-deck structure is observed when localized surface perturbations exist. For demonstration, we focus on the steady triple-deck flow over an incompressible two-dimensional (2D) flat plate with a localized roughness element. The roughness is located at a distance $L$ downstream from the leading edge. Taking $L$ as the reference length, the Reynolds number is defined as $\text{Re}=\frac{U_{\infty}L}{\nu}$, with $U_{\infty}$ and $\nu$ being the velocity of the incoming flow and the kinematic viscosity, respectively. For convenience, we introduce a small parameter $\epsilon=\text{Re}^{-\frac{1}{8}}\ll 1$. The system is described in the Cartesian coordinate $(X,Y)$, with its origin at the leading edge. The roughness height is assumed to be $O(\epsilon^{5})$ or smaller, and we denote it as $\{(X,Y)\big{|}\ Y=\epsilon^{5}F(\frac{X-1}{\epsilon^{3}}),\ X\in{\mathbb{R}}\}$ for some given function $F=O(1)$ as $\epsilon\to 0$. Consequently, the roughness-induced mean-flow distortion exhibits three distinct layers in the transverse direction, spanning in the $O(\epsilon^{3})$ vicinity of the roughness (see figure 1).
@@
We denote the velocity vector by ${\mathbf{u}}_{NS}=(u_{NS}(X,Y),v_{NS}(X,Y))$, the pressure $p_{NS}$, and demonstrate their scales in Table 1.
@@ I II III
$u_{NS}$ $O(\epsilon)$ $\tilde{U}_{0}(\frac{Y}{\epsilon^{4}})+O(\epsilon)$ $1+O(\epsilon^{2})$
$v_{NS}$ $O(\epsilon^{3})$ $O(\epsilon^{2})$ $O(\epsilon^{2})$
$p_{NS}$ $O(\epsilon^{2})$ $O(\epsilon^{2})$ $O(\epsilon^{2})$
@@
Based on the scaling relations in Figure 1 and Table 1, and by applying the method of matched asymptotic expansions (see details in Appendix), we derive that the flows in the upper layer and main layer are given by
(1.2) $u_{NS}\sim 1-\epsilon\Re(f(z)),\quad v_{NS}\sim\epsilon^{2}\Im(f(z)),\quad p_{NS}\sim\epsilon^{2}\Re(f(z)),\quad z=x_{*}+i\bar{y},$
$u_{NS}\sim\tilde{U}_{0}(\tilde{y})+\epsilon A_{*}(x_{*})\tilde{U}_{0}^{\prime}(\tilde{y}),\quad v_{NS}\sim-\epsilon^{2}A_{*}^{\prime}(x_{*})\tilde{U}_{0}(\tilde{y}),\quad p_{NS}\sim\epsilon^{2}\tilde{p}(x_{*}).$
Here $x_{*}=\frac{X-1}{\epsilon^{3}},\ \bar{y}=\frac{Y}{\epsilon^{3}},\ \tilde{y}=\frac{Y}{\epsilon^{4}}$, $\tilde{U}_{0}^{\prime}(0)=1$, $A_{*}$ is an unknown function (called the displacement of the velocity), and
(1.3) $f(z)=-\frac{1}{\pi}\int_{{\mathbb{R}}}\frac{A_{*}^{\prime}}{\zeta-z}d\zeta,\quad\tilde{p}(x_{*})=\lim_{\Im(z)\to 0^{+}}\Re(f(z))=-\frac{1}{\pi}\text{P.V.}\int_{{\mathbb{R}}}\frac{A_{*}^{\prime}(\zeta)}{\zeta-x_{*}}d\zeta.$
For the viscous sublayer (region I), we derive the governing equations as
(1.4) $\left\{\begin{aligned} &u_{*}\partial_{x_{*}}u_{*}+v_{*}\partial_{y_{*}}u_{*}+\frac{dP_{*}}{dx_{*}}-\partial_{y_{*}}^{2}u_{*}=0,\\ &\partial_{x_{*}}u_{*}+\partial_{y_{*}}v_{*}=0,\end{aligned}\right$.
@@
To close the system, the upper boundary condition of the sublayer equations is enforced by matching with the main layer solution. Now, we consider the overlapping region connecting the sublayer and main layer, where $y_{*}=O(\epsilon^{-\alpha_{0}})$ with $\alpha_{0}\in(0,1)$. Since the Blasius solution satisfies $\tilde{U}_{0}\to\tilde{y}+O(\tilde{y}^{4})$ as $\tilde{y}\to 0^{+}$, then for any $\beta_{0}>0$, it follows that
(1.7) $u_{NS}=$ $\ \tilde{U}_{0}(\tilde{y})+\epsilon A_{*}(x_{*})\tilde{U}_{0}^{\prime}(\tilde{y})+o(\epsilon)=\tilde{y}+\epsilon A_{*}(x_{*})+o(\epsilon+\tilde{y}^{4-\beta_{0}})$
$=$ $\ \epsilon\big{(}y_{*}+A_{*}(x_{*})\big{)}+o\big{(}\epsilon+\epsilon^{(4-\beta_{0})(1-\alpha_{0})}\big{)}.$
To ensure the second part to be negligible, we require $(4-\beta_{0})(1-\alpha_{0})\geq 1$, namely, $\alpha_{0}\in(0,\frac{3}{4})$ and $\beta_{0}\in(0,4)$. Likewise, we can derive that the leading-order pressure in the sublayer and that in the main layer are equal, that is,
@@
In light of (1.7), it is permissible to set the upper boundary of the viscous sublyaer domain to be a curve $\gamma=\{(x_{*},y_{*})|\ y_{*}=f(x_{*})\in O(\epsilon^{-\alpha_{0}}),\ x_{*}\in{\mathbb{R}}\}$, where $f$ is to be determined. Consequently, the upper boundary condition is
@@
By (1.4)-(1.6), and (1.9), we derive the closed system for the viscous sublayer as
@@
A critical observation is that the flows (1.2) in the upper two layers only depend on the unknown quantities $(u_{*},v_{*},A_{*},P_{*})$ within the viscous sublayer, this paper thus focuses on the equations (1.10) of the viscous sublayer (henceforth referred to as the triple-deck equations).
1.1. Known results
Mathematically, the triple-deck equations share a structural resemblance to the Prandtl equations of classical boundary layers, yet they differ fundamentally in domain and pressure determination. The Prandtl equations are defined over $(0,L_{0})\times\mathbb{R}_{+}$, with pressure $P_{*}$ dictated by external potential flow $u_{E}(x_{*})$ via the Bernoulli relation
@@ 1.1. Known results
which means that the pressure is a given function. There are a lot of results on this topic. The global existence of solutions to the Prandtl equations was first established by Oleinik [19], and subsequent decades have yielded rich results on well-posedness, asymptotic behavior, and regularity [7, 8, 16, 23, 24, 29, 30]. Moreover, the flow separation and reverse phenomena have also attracted the attention of many scholars [2, 10, 15, 25].
Compared with (1.11), the pressure of triple-deck model is given by (1.8) which is defined by unknown function $A_{*}$. Due to that, to construct a solution of triple-deck model is not an easy way. For the unsteady case, Iyer and Vicol [11], Gérard-Varet et al. [5] proved the well-posedness of solutions in real analytic spaces, while Dietert and Gérard-Varet [3] revealed instability for some smooth initial data, indicating that the class of analytic functions is optimal for well-posedness. For steady cases, mathematical results are sparser. When $F=0$, Iyer and Maekawa [9] proved a local Liouville-type theorem, pointing out that under specific rigid conditions, the equations admit only the Couette solution. However, physicists found that the triple-deck structure only appears in flow with some perturbations which means that $F\not\equiv 0$. For more results on physical analysis and numerical simulations, see [1, 4, 12, 26, 32]. It is noteworthy that, for the triple-deck equations (1.10) in the presence of disturbance, the well-posedness of solutions lacks rigorous mathematical conclusions. This paper investigates the triple-deck equations (1.10) for some $f=O(\epsilon^{-\alpha_{0}})$ with $\alpha_{0}\in(0,\frac{3}{4})$, establishes the existence and uniqueness theory in the presence of disturbance for the first time, and verifies the compatibility between the disturbance scale and physical observation results.
@@ 1.2. Main results
To simplify the equations (1.10), we introduce the new coordinate $(x,y)$ and new notations $(u,v,P,A)$:
@@
Then, the problem (1.10) can be reduced to
@@
For convenience, we take $f(x)=M+F(x)$ for $M=\epsilon^{-\alpha_{0}}$ and $\alpha_{0}\in(0,\frac{3}{4})$. It is useful to introduce the vorticity $\omega=\partial_{y}u$ and
@@
Differentiating the first equation of (1.13) in the variable $y$, we obtain the perturbation system about $(\omega,A)$ as follows.
@@
Here $F$ is a prescribed wall disturbance function. We denote $\Omega_{M}={\mathbb{R}}\times[0,M]$.
To establish a rigorous framework, we introduce some functional spaces in this paper. Define the weighted Sobolev spaces for vorticity $\omega$ as
(1.16) $X_{\bar{\alpha}}=$ $\ \bigg{\{}f\big{|}\ {\big{\|}f\big{\|}}_{X_{\bar{\alpha}}}={\big{\|}f\big{\|}}_{L^{2}(\Omega_{M})}+{\big{\|}|\partial_{x}|^{\frac{1}{18}}y^{\frac{1}{6}}f\big{\|}}_{L^{2}(\Omega_{M})}+{\big{\|}f\big{\|}}_{Y_{\bar{\alpha}}}<\infty\bigg{\}},$
@@
where $|\partial_{x}|^{s}:=\mathcal{F}^{-1}[|\xi|^{s}\mathcal{F}[\cdot]]$ denotes the Fourier multiplier of order $s$.
For the displacement function $A$, we define
(1.17) $X_{\alpha,\infty}=\bigg{\{}f\big{|}\ {\big{\|}|\partial_{x}|^{\frac{5}{6}}f\big{\|}}_{H^{\frac{4}{3}+\frac{\alpha}{2}}({\mathbb{R}})}+{\big{\|}f\big{\|}}_{L^{\infty}({\mathbb{R}})}<\infty\bigg{\}}.$
The combined space for $(\omega,A)$ incorporates the upper boundary condition is
(1.18) ${\mathcal{X}}_{\alpha,M}=\bigg{\{}(\omega,A)\big{|}$ $\ {\big{\|}(\omega,A)\big{\|}}_{{\mathcal{X}}_{\alpha,M}}={\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}A\big{\|}}_{X_{\alpha,\infty}}<\infty,$
@@
We adopt standard mixed-norm notations on $\Omega_{M}$ as following:
(1.19) $\langle f,g\rangle_{L^{2}(\Omega_{M})}=\int_{\Omega_{M}}f\bar{g}\ d{\mathbf{x}},\quad{\big{\|}f\big{\|}}_{L^{2}}^{2}=\int_{\Omega_{M}}|f|^{2}d{\mathbf{x}},$
For some inequalities, the notation $B_{0}\lesssim B_{1}$ denotes that there exists some $C>0$ independent of $B_{0}$ and $B_{1}$, such that $B_{0}\leq CB_{1}$.
To seek the solution of the system (1.15), we use the following decomposition $(\omega,A)=(\omega_{0}+\bar{\omega},A_{0}+\bar{A})$ where $(\omega_{0},A_{0})$ satisfies the following linear part of (1.15):
(1.20) $\left\{\begin{aligned} &y\partial_{x}\omega_{0}-\partial_{y}^{2}\omega_{0}=0,\quad(x,y)\in{\mathbb{R}}\times[0,M],\\ &\partial_{y}\omega_{0}\big{|}_{y=0}=\partial_{x}|\partial_{x}|A_{0},\quad I_{M}[\omega_{0}]=A_{0}(x)+F(x),\quad x\in{\mathbb{R}}.\end{aligned}\right$.
Our first result is to construct a solution of (1.20).
Theorem 1.1.
Let $\alpha\in(0,\frac{7}{3}]$, then the linearized system (1.20) has a solutions $(\omega_{0},A_{0})$ belongs to ${\mathcal{X}}_{\alpha,M}$ and satisfies the following estimates:
where $C>0$ is independent of $\alpha,M$, and $F$.
Now, we are the position to state the main results of this paper.
Theorem 1.2.
Let $\alpha\in(\frac{2}{3},\frac{7}{3}]$ and $\varepsilon\in(0,\frac{1}{6}]$. Then, there exist a small enough constant $\delta_{0}=\delta_{0}(\alpha,\varepsilon)>0$ and large enough constant $M_{0}$, such that when $M\geq M_{0}$ and
the system (1.15) admits a solution $(\omega,A)=(\omega_{0}+\bar{\omega},A_{0}+\bar{A})\in{\mathcal{X}}_{\alpha,M}$ where $(\omega_{0},A_{0})$ defined by (1.20) and $(\bar{\omega},\bar{A})$ satisfying
(1.22) ${\big{\|}\bar{\omega}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\bar{A}\big{\|}}_{X_{\alpha,\infty}}\leq$ $\ CM^{\frac{5+3\varepsilon}{2}}{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}^{2},$
where $C>0$ is independent of $\alpha,\varepsilon,M$, and $F$. Moreover, $(\omega,A)$ is the unique solution of (1.15) within ${\mathcal{X}}_{\alpha,M}$.
Remark 1.1.
Our theoretical results align closely with physical phenomena. The linear component $(\omega_{0},A_{0})$ captures the dominant behavior of $(\omega,A)$ under small perturbations $F$, consistent with the principle that linear approximations govern near-equilibrium states in fluid dynamics.
Remark 1.2.
Considering the scaling relation $M=O(\text{Re}^{\frac{\alpha_{0}}{8}})$ for $\alpha_{0}\in(0,\frac{3}{4})$, our analysis shows that the allowable perturbation scale $O(\text{Re}^{-a_{*}})$ with $a_{*}\in(0,\infty)$ falls within the physically reasonable range $0,\infty)$ (see [26]).
Remark 1.3.
When $F=0$, our results imply that there is a unique zero solution in ${\mathcal{X}}_{\alpha,M}$ which aligns with the results by Iyer and Maekawa in [9].
Remark 1.4.
Here, we take $M\sim\epsilon^{-\alpha_{0}}$ with some small data $\epsilon$ which implies that $M$ will be large enough. Thus, taking $M_{0}$ large enough is satisfied by our model. On the other hand, for general $M_{0}$, the results may also be got by Remark 3.2.
1.3. Main ideas
In this subsection, we state the main ideas of the paper. First, we notice that the system (1.15) is a degenerate nonlinear parabolic equation with variable coefficients, where the degeneration occurs on the wall $y=0$. To deal with the degeneration, we decompose $\omega=\omega_{0}+\bar{\omega}$ and $A=A_{0}+\bar{A}$ satisfying
@@ 1.3. Main ideas
It separates the main effects caused by the roughness $F$ from the mathematical difficulties arising from nonlinear terms.
In the homogeneous problem (1.23), we observe that the integral condition exhibits nonlocal behavior, so it cannot be treated in isolation. Instead, a unified framework is needed to couple this integral condition with the boundary data at $y=0$ (i.e., $\partial_{y}\omega_{0}\big{|}_{y=0}=\partial_{x}|\partial_{x}|A_{0}$). Their synthesis yields an elliptic multiplier $m(D)^{-1}$ (see (3.45)). This multiplier crucially links the displacement $A_{0}$ and vorticity $\omega_{0}$ by
(1.25) ${\big{\|}\omega_{0}\big{\|}}_{X_{\bar{\alpha}}}\lesssim{\big{\|}m(D)^{-1}|\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}A_{0}\big{\|}}_{L^{2}}\lesssim{\big{\|}F\big{\|}}_{\dot{H}^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}({\mathbb{R}})},\ \bar{\alpha}\geq 0,$
providing a key tool for analyzing the well-posedness of the pair $(\omega,A)$.
In order to solve the non-homogeneous problem (1.24), the nonlinear terms and boundary terms are treated separately. Then, we decompose $\bar{\omega}=\bar{\omega}_{e}+\bar{\omega}_{b}$ satisfying
(1.26) $\left\{\begin{aligned} &y\partial_{x}\bar{\omega}_{e}-\partial_{y}^{2}\bar{\omega}_{e}=-(u\partial_{x}\omega+v\partial_{y}\omega),\quad(x,y)\in\Omega_{M},\\ &u=I_{y}[\omega],\quad v=-\partial_{x}I_{y}[u],\quad(x,y)\in\Omega_{M}\\ &\partial_{y}\bar{\omega}_{e}\big{|}_{y=0}=0,\quad x\in{\mathbb{R}},\end{aligned}\right$.
and
(1.27) $\left\{\begin{aligned} &y\partial_{x}\bar{\omega}_{b}-\partial_{y}^{2}\bar{\omega}_{b}=0,\quad(x,y)\in\Omega_{M}\\ &\partial_{y}\bar{\omega}_{b}\big{|}_{y=0}=\partial_{x}|\partial_{x}|\bar{A},\quad I_{M}[\bar{\omega}_{b}]=\bar{A}-I_{M}[\bar{\omega}_{e}],\quad x\in{\mathbb{R}}.\end{aligned}\right$.
Similar to the homogeneous problem (1.23), we can find a solution $(\bar{\omega}_{b},\bar{A})$ to (1.27) with the following estimates
(1.28) ${\big{\|}\bar{\omega}_{b}\big{\|}}_{X_{\bar{\alpha}}}\lesssim$ $\ {\big{\|}m(D)^{-1}|\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}\bar{A}\big{\|}}_{L^{2}}\lesssim{\big{\|}I_{M}[\bar{\omega}_{e}]\big{\|}}_{\dot{H}^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}({\mathbb{R}})}\quad\text{for }\ \bar{\alpha}\geq 0$
[…]
Hence, the essential step is to solve (1.26). To this end, we firstly omit the integral condition and consider a linear problem as following

### v2

# On the classical solution for the steady triple-deck equations

@@ Abstract.
This paper establishes the existence and uniqueness of classical solutions to the steady Triple-Deck equations, which describe incompressible boundary layer flow over localized roughness at high Reynolds numbers. The triple-deck theory was developed to overcome the Goldstein singularity in classical Prandtl boundary layer theory, capturing the interaction between the viscous sublayer, main layer, and upper layer when surface roughness of height $O({\rm Re}^{-\frac{5}{8}})$ is present. The key ingredients of this paper include: (1) A decomposition separating roughness effects from nonlinear difficulties; (2) A novel Green’s function using Airy functions that overcomes low-frequency singularities via the non-vanishing property of $\sqrt{3}Ai(z)+Bi(z)$; (3) The introduction of weighted Sobolev norms ${\big\||\partial_{x}|^{\frac{1}{18}}y^{\frac{1}{6}}\omega\big\|}_{L^{2}}$ of the vorticity yielding $M$-independent estimates for displacement $A$. As a byproduct, local uniqueness of Couette flow is established when $F=0$.
@@ 1. Introduction
1.1. The Triple-Deck equations
In 1904, Prandtl [21] proposed the boundary layer theory to address attached flow problems, achieving remarkable success in fluid dynamics research. However, when dealing with flow separation induced by adverse pressure gradients, this theory encounters singularity issues. Notably, Landau and Lifshitz [14] revealed that the velocity component normal to the body surface grows unboundedly (termed the Goldstein singularity), which hinders the solution from being further extended to the reverse-flow region [6]. To overcome these difficulties, scholars including Messiter [17] and Stewartson [27] developed the triple-deck theory from the late 1960s to the early 1970s. In this theory, three distinct layers—a viscous sublayer, a main layer, and an upper layer—appear in the transverse direction, for which uniformly valid approximate solutions are derived by the asymptotic matched method.
The 2D steady incompressible Navier-Stokes equations take the following dimensionless form:
Here $(u_{NS},v_{NS})$ denotes the velocity vector, $p_{NS}$ pressure, and $\epsilon={\rm Re}^{\frac{1}{8}}$ with the Reynolds number ${\rm Re}\gg 1$. In physics, the triple-deck structure emerges when localized surface roughness of height ${\mathcal{O}}(\epsilon^{5})$ is present. We then denote it as
$\Big\{(X,Y)\big|\ Y=\epsilon^{5}F\big(\frac{X-1}{\epsilon^{3}}\big),\ X\in{\mathbb{R}}\Big\}$
for some given function $F={\mathcal{O}}(1)$ as $\epsilon\to 0$. Consequently, the roughness-induced mean-flow distortion exhibits three distinct layers in the transverse direction, spanning in the ${\mathcal{O}}(\epsilon^{3})$ vicinity of the roughness (see figure 1).
@@ 1.1. The Triple-Deck equations
We demonstrate the scales of $(u_{NS},v_{NS},p_{NS})$ in Table 1.
@@ I II III
$u_{NS}$ ${\mathcal{O}}(\epsilon)$ $\tilde{U}_{0}(\frac{Y}{\epsilon^{4}})+{\mathcal{O}}(\epsilon)$ $1+{\mathcal{O}}(\epsilon^{2})$
$v_{NS}$ ${\mathcal{O}}(\epsilon^{3})$ ${\mathcal{O}}(\epsilon^{2})$ ${\mathcal{O}}(\epsilon^{2})$
$p_{NS}$ ${\mathcal{O}}(\epsilon^{2})$ ${\mathcal{O}}(\epsilon^{2})$ ${\mathcal{O}}(\epsilon^{2})$
@@
Based on the scaling relations in Figure 1 and Table 1, and via the method of matched asymptotic expansions applied to (1.1) (see Appendix for details), we derive the governing equations for the viscous sublayer (region I) as
(1.2) $\left\{\begin{aligned} &u_{*}\partial_{x_{*}}u_{*}+v_{*}\partial_{y_{*}}u_{*}+\frac{dP_{*}}{dx_{*}}-\partial_{y_{*}}^{2}u_{*}=0,\\ &\partial_{x_{*}}u_{*}+\partial_{y_{*}}v_{*}=0.\end{aligned}\right$.
@@
To close the system, the upper boundary condition of the sublayer equations is enforced by matching with the main layer solution (see Appendix for details). Then we can derive the leading-order pressure as follows
@@
And the upper boundary condition of $u_{*}$ is
@@
Here $f$ is given (to be determined later). Consequently, by (1.2)-(1.6), we derive the closed system for the viscous sublayer as
@@
The system (1.7) is referred to as the Triple-Deck equations.
Recall that the Prandtl equations of classical boundary layers take
(1.8) $\left\{\begin{aligned} &u_{*}\partial_{x_{*}}u_{*}+v_{*}\partial_{y_{*}}u_{*}+\frac{dP_{*}}{dx_{*}}-\partial_{y_{*}}^{2}u_{*}=0,\quad\text{in }(0,L_{0})\times{\mathbb{R}}_{+},\\ &\partial_{x_{*}}u_{*}+\partial_{y_{*}}v_{*}=0,\quad\text{in }(0,L_{0})\times{\mathbb{R}}_{+},\\ &(u_{*},v_{*})\big|_{y_{*}=0}=0,\quad x_{*}\in{\mathbb{R}}_{+},\\ &u_{*}\big|_{x_{*}=0}=u_{0},\quad y_{*}\in{\mathbb{R}}_{+},\end{aligned}\right$.
where $u_{0}$ is prescribed. Mathematically, the Triple-Deck equations share a structural resemblance to the Prandtl equations (1.8), yet they differ fundamentally in domain and pressure determination. The Prandtl equations are defined over $(0,L_{0})\times\mathbb{R}_{+}$, with pressure $P_{*}$ dictated by external potential flow $u_{E}(x_{*})$ via the Bernoulli relation
@@
which means that the pressure is a given function. While, the pressure of Triple-Deck model is given by (1.5) with the unknown function $A_{*}$.
The existence and uniqueness of classical solutions to the Prandtl equations were first established by Oleinik [19]; subsequent decades have yielded rich results on well-posedness, asymptotic behavior, and regularity [7, 8, 16, 23, 24, 29, 30]. Moreover, flow separation and reverse-flow phenomena have attracted considerable attention from many scholars [2, 10, 15, 25].
When $F=0$, it is easy to see that $(u_{*},v_{*},P_{*},A_{*})=(y,0,0,0)$ is a trivial solution to (1.7). In a recent work [9], Iyer and Maekawa proved the rigidity of this trivial solution under certain assumptions on the solution. However, physicists found that the triple-deck structure only appears when $F\not\equiv 0$. Since the $A_{*}$ remains unknown in this case, finding a non-trivial solution becomes a highly non-trivial problem. Numerous results are available on physical analysis and numerical simulations; see [1, 4, 12, 26, 32]. Yet, if $F\not\equiv 0$, the well-posedness of (1.7) still lacks rigorous mathematical justification. For the unsteady case, Iyer and Vicol [11] and Gérard-Varet et al. [5] established well-posedness in real-analytic spaces, whereas Dietert and Gérard-Varet [3] exhibited instability for some smooth initial data, indicating that the analytic class is optimal for well-posedness.
@@ 1.2. Main results
To simplify the system (1.7), we introduce the new coordinate $(x,y)$ and new unknowns $(u,v,P,A)$:
@@
Then the problem (1.7) can be reduced to
@@
For convenience, we take $f(x)=M+F(x)$ for $M>0$. We introduce the vorticity $\omega=\partial_{y}u$ and
@@
Differentiating the first equation of (1.11) in the variable $y$, we obtain the perturbation system about $(\omega,A)$ as follows
@@
Here $F$ is a prescribed function. We denote $\Omega_{M}={\mathbb{R}}\times[0,M]$. It is easy to see that the Couette flow $(u_{*},v_{*},P_{*},A_{*})$ corresponds to the trivial solution $(u,v,P,A)=(0,0,0,0)$ when $F=0$.
Now we introduce the functional framework of classical solutions. For the vorticity $\omega$, we define the weighted Sobolev spaces by
$X_{\bar{\alpha}}=$ $\ \bigg\{f\big|\ {\big\|f\big\|}_{X_{\bar{\alpha}}}={\big\|f\big\|}_{L^{2}(\Omega_{M})}+{\big\||\partial_{x}|^{\frac{1}{18}}y^{\frac{1}{6}}f\big\|}_{L^{2}(\Omega_{M})}+{\big\|f\big\|}_{Y_{\bar{\alpha}}}<\infty\bigg\},$
@@
where the parameter $\bar{\alpha}$ is nonnegative and $|\partial_{x}|^{s}:=\mathcal{F}^{-1}[|\xi|^{s}\mathcal{F}[\cdot]]$ denotes the Fourier multiplier of order $s$. For the displacement function $A$, we define
$X_{\alpha,\infty}=\bigg\{f\big|\ {\big\||\partial_{x}|^{\frac{5}{6}}f\big\|}_{H^{\frac{4}{3}+\frac{\alpha}{2}}({\mathbb{R}})}+{\big\|f\big\|}_{L^{\infty}({\mathbb{R}})}<\infty\bigg\}.$
Then the functional space for $(\omega,A)$ incorporating the upper boundary condition is defined by
@@
To seek the solution to the system (1.13), we introduce the following crucial decomposition $(\omega,A)=(\omega_{0}+\bar{\omega},A_{0}+\bar{A})$, where $(\omega_{0},A_{0})$ and $(\bar{\omega},\bar{A})$ solve the following system
@@
Our first main result is the solvability of the linear system (1.14).
Theorem 1.1.
Let $\alpha\in(0,\frac{7}{3}]$. The linear system (1.14) admits a solution $(\omega_{0},A_{0})\in{\mathcal{X}}_{\alpha,M}$ satisfying the estimate
where $C>0$ is independent of $\alpha,M$.
The second main result of this paper is the existence and uniqueness of classical solution to the nonlinear system (1.13).
Theorem 1.2.
Let $\alpha\in(\frac{2}{3},\frac{7}{3}]$ and $\varepsilon\in(0,\frac{1}{6}]$. There exist a small constant $\delta_{0}=\delta_{0}(\alpha,\varepsilon)>0$ and a sufficiently large constant $M_{0}$, such that when $M\geq M_{0}$ and
the system (1.13) admits a solution $(\omega,A)=(\omega_{0}+\bar{\omega},A_{0}+\bar{A})\in{\mathcal{X}}_{\alpha,M}$ with
${\big\|\bar{\omega}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\bar{A}\big\|}_{X_{\alpha,\infty}}\leq$ $\ CM^{\frac{5+3\varepsilon}{2}}{\big\|F\big\|}_{H^{2}({\mathbb{R}})}^{2}.$
Here $C>0$ is independent of $\alpha,\varepsilon,M$. Moreover, $(\omega,A)$ is the unique solution to the system (1.13) within ${\mathcal{X}}_{\alpha,M}$ if
${\big\|(\omega,A)\big\|}_{{\mathcal{X}}_{\alpha,M}}+{\big\|F\big\|}_{H^{2}({\mathbb{R}})}\leq 2\delta_{0}M^{-\frac{5+3\varepsilon}{2}}.$
Remark 1.1.
Our theoretical results align closely with the physical phenomena. The linear component $(\omega_{0},A_{0})$ captures the dominant behavior of $(\omega,A)$. Considering the scaling relation $M={\mathcal{O}}(\text{Re}^{\frac{\alpha_{0}}{8}})$ for $\alpha_{0}\in(0,\frac{3}{4})$ (see Appendix for details), our analysis shows that the allowable roughness scale ${\mathcal{O}}(\text{Re}^{-a_{*}})$ with $a_{*}\in(0,\infty)$ falls within the physically reasonable range $0,\infty)$ (see [26]).
Remark 1.2.
If $F=0$, Iyer and Maekawa [9] derived the local rigidity of Couette flow. Specifically, they proved that an $L^{2}$ strong solution $(\omega,A)$ of (1.13) is trivial if ${\big\|(\omega,A)\big\|}_{sc}$ is sufficiently small, where ${\big\|\cdot\big\|}_{sc}$ is a scaling invariant norm. As a byproduct of Theorem 1.2, we further establish the local uniqueness of Couette flow: the system (1.13) has a locally unique solution in ${\mathcal{X}}_{\alpha,M}$, which is a trivial solution.
Remark 1.3.
We adopt $M\gg 1$ in our model for consistency with physical observations. Indeed, for general $M>0$, the results may also be got by Remark 3.2.
1.3. Main difficulties and Key ideas
The system (1.13) is a degenerate nonlinear parabolic equation with variable coefficients and its degeneration occurs on the wall $y=0$. To solve the system (1.13), we first introduce a crucial decomposition (1.14) and (1.15), which separates the confluence caused by the roughness $F$ from the mathematical difficulties arising from nonlinear terms.
For the homogeneous problem (1.14), the key ingredient is to solve the following linear problem